Find the mean, median and mode for each set of data:
a) 13.5, 15, 13.5,11,13
b) 32, 28.3, 26.8, 31, 24.4

Answer by Mathtut(3670)

a) mean (13.5+15+13.5+11+13)/5(terms)=66/5=13.2 is the mean
:
median of {11,13,13.5,13.5,15} is 13.5 the middle value
:
mode is also 13.5 as it is the number repeated most often
:
b) 32, 28.3, 26.8, 31, 24.4
re write in ascending order to {24.4,26.8,28.3,31,32}
:
mean 24.4+26.8+28.3+31+32/5(terms)=142.5/5=28.5
:
median is 28.3 the middle term
:
mode is: If all the numbers appear only once, then there is no mode. Therefore there is no mode in this problem